i need to solve this equation from standard form to vertex form with all the solutions please and thank you
h(t)= -4.9t^2+2t+45

Answer by stanbon(75887)

i need to solve this equation from standard form to vertex form with all the solutions please and thank you
h(t)= -4.9t^2+2t+45
-----
4.9t^2 -2t -45 = 0
4.9(t^2 - (2/4.9)t + ?) = 45 + 4.9*?
----
4.9(t^2 - (2/4.9)t + (1/4.9)^2 = 45 + 4.9(1/4.9)^2
----
4.9(t-1/4.9)^2 = 45.204
-----
Vertex( 1/4.9 , 45.204)
(Cheers,
Stan H.
=========